OpenStudy (anonymous):

f(x)=1+3x-x^2 find the min and max values of the function

OpenStudy (anonymous):

this is a parabola that faces down, so it has no minimum maximum is the second coordinate of the vertex first coordinate of the vertex is always \(-\frac{b}{2a}\) in your case it is \(-\frac{3}{2\times (-1)}=\frac{3}{2}\) replace \(x\) by \(\frac{3}{2}\) to get the maximum

OpenStudy (anonymous):

isnt that 3/-2??

OpenStudy (calculusfunctions):

Do you want a calculus based or a non-calculus based solution?

OpenStudy (anonymous):

non calc based

OpenStudy (calculusfunctions):

Alright! First rearrange the original polynomial to the form f(x) = ax² + bx + c Hence f(x) = -x² + 3x + 1 Now complete the square f(x) = (-x² + 3x) + 1 [step 1: separate the variable terms from the constant] f(x) = -(x² - 3x) + 1 [step 2: factor the coefficient of x² from the parentheses] f(x) = -(x² - 3x + 9/4 - 9/4) + 1 {step 3:add and subtract the square of half the coefficient of x inside the parentheses] f(x) = -(x² - 3x + 9/4) + (-9/4)(-1) + 1 [step 4: remove the -9/4 from the parentheses and multiply by the original coefficient of x² that was factored in step 2] f(x) = -(x - 3/2)² + 13/4 [step 5: a² - 2ab + b² = (a - b)²] The vertex of this parabola is at (3/2, 13/4) and the parabola's direction of opening is downward. ∴ the maximum value is 13/4 and occurs at x = 3/2. Note: It is not necessary to perform all 5 steps of completing the square, however when I'm teaching, I like explaining everything properly. I really hope that helps!

OpenStudy (anonymous):

no it is \(\frac{3}{2}\) because there is a minus sign in the formula, and also \(a=-1\)

OpenStudy (calculusfunctions):

The maximum value is NOT 3/2. An optimum value is the y-coordinate of the vertex and occurs at the x-coordinate of the vertex. THEREFORE THE MAXIMUM VALUE IS 13/4 WHEN x = 3/2.